The soul package

Melchior FRANZ

November 17, 2003

Abstract

This article describes the soul package^{1}, which provides h y p h e n - a t a b l e l e t t e r s p a c i n g ( s p a c i n g o u t ) , underlining and some derivatives such as overstriking and highlighting. Although the package is optimized for L^{A}TEX 2ε, it also works with Plain TEX and with other flavors of TEX like, for instance, ConTEXt. By the way, the package name soul is only a combination of the two macro names \so (space out) and \ul (underline)—nothing poetic at all.

Contents

1 Typesetting rules 2

2 Short introduction and common rules 3 2.1 Some things work . . . . . 3 2.2 … others don’t . . . . . . 5 2.3 Troubleshooting . . . . . . 6

3 L e t t e r s p a c i n g 8 3.1 How it works . . . . . . . 8 3.2 Some examples . . . . . . 8 3.3 Typesetting caps-and-

small-caps fonts …. 10 3.4 Typesetting Fraktur … 11 3.5 Dirty tricks . . . . . . . . 11

4 Underlining 11 4.1 Settings . . . . . . . . . . 12 4.2 Some examples . . . . . . 13

5 Customization 14 5.1 Adding accents . . . . . . 14 5.2 Adding font commands . 15 5.3 Changing the internal font 15 5.4 The configuration file … 15 6 Miscellaneous 16 6.1 Using soul with other

flavors of TEX . . . . . . . 16 6.2 Using soul commands

for logical markup …. 17 6.3 Typesetting long words

in narrow columns …. 19 6.4 Using soul commands in

section headings . . . . . 19

7 How the package works 21 7.1 The kernel . . . . . . . . . 21 7.2 The interface . . . . . . . 22 7.3 A driver example . . . . . 24

8 The implementation 26 8.1 The kernel . . . . . . . . . 28 8.2 The scanner . . . . . . . . 29 8.3 The analyzer . . . . . . . 34 8.4 The l e t t e r s p a c i n g

driver . . . . . . . . . . . 38 8.5 The caps driver . . . . . 42 8.6 The underlining driver . . 44 8.7 The overstriking driver . . 46 8.8 The highlighting driver . . 47

1This file has version number 2.4, last revised 2003/11/17.

1 Typesetting rules

There are several possibilities to emphasize parts of a paragraph, not all of which are considered good style. While underlining is commonly rejected, experts dispute about whether letterspacing should be used or not, and in which cases. If you are not interested in such debates, you may well skip to the next section.

Theory …

To understand the experts’ arguments we have to know about the conception of page grayness. The sum of all characters on a page represents a certain amount of grayness, provided that the letters are printed black onto white paper.

Jan Tschichold [10], a well known and recognized typographer, accepts only forms of emphasizing, which do not disturb this grayness. This is only true of italic shape, caps, and caps-and-small-caps fonts, but not of ordinary letterspacing, underlining, bold face type and so on, all of which appear as either dark or light spots in the text area. In his opinion emphasized text shall not catch the eye when running over the text, but rather when actually reading the respective words.

Other, less restrictive typographers [11] call this kind of emphasizing ‘integrated’ or ‘aesthetic’, while they describe ‘active’ emphasizing apart from it, which actually has to catch the reader’s eye. To the latter group belong commonly despised things like letterspacing, demibold face type and even underlined and colored text.

On the other hand, Tschichold suggests to space out caps and caps-and-small-caps fonts on title pages, headings and running headers from 1 pt up to 2 pt. Even in running text legibility of uppercase letters should be improved with slight letterspacing, since (the Roman) majuscules don’t look right, if they are spaced like (the Carolingian) minuscules.^{2}

… and Practice

However, in the last centuries letterspacing was excessively used, underlining at least sometimes, because capitals and italic shape could not be used together with the Fraktur font and other black-letter fonts, which are sometimes also called “old German” fonts. This tradition is widely continued until today. The same limitations apply still today to many languages with non-latin glyphs, which is why letterspacing has a strong tradition in eastern countries where Cyrillic fonts are used.

The Duden [4], a well known German dictionary, explains how to space out properly: Punctuation marks are spaced out like letters, except quotation marks and periods. Numbers are never spaced out. The German syllable -sche is not spaced out in cases like “der V i r c h o w sche Versuch”^{3}. In the old German Fraktur fonts the ligatures ch, ck, sz (ß) and tz are not broken within spaced out text.

While some books follow all these rules [6], others don’t [7]. In fact, most books in my personal library do not space out commas.

2This suggestion is followed throughout this article, although Prof. Knuth already considered slight letterspacing with his cmcsc fonts.

3the Virchow experiment

2 Short introduction and common rules

The soul package provides five commands that are aimed at emphasizing text parts. Each of the commands takes one argument that can either be the text itself or the name of a macro that contains text (e. g. \so\text)^{4}. See table 1 for a complete command survey.

                \so{letterspacing} l e t t e r s p a c i n g
   \caps{CAPITALS, Small Capitals} CAPITALS, Small Capitals
                  \ul{underlining} underlining
                 \st{overstriking} overstriking
                 \hl{highlighting} highlighting^{5}

The \hl command does only highlight if the color package was loaded, otherwise it falls back to underlining.^{6} The highlighting color is by default yellow, underlines and overstriking lines are by default black. The colors can be changed using the following commands:

\setulcolor{red} set underlining color
\setstcolor{green} set overstriking color
\sethlcolor{blue} set highlighting color

\setulcolor{} and \setstcolor{} turn coloring off. There are only few colors predefined by the color package, but you can easily add custom color definitions. See the color package documentation [3] for further information.

    \usepackage{color,soul}
    \definecolor{lightblue}{rgb}{.90,.95,1}
    \sethlcolor{lightblue}
    ...
    \hl{this is highlighted in light blue color}

2.1 Some things work …

The following examples may look boring and redundant, because they describe nothing else than common L^{A}TEX notation with a few exceptions, but this is only the half story: The soul package has to pre-process the argument before it can split it into characters and syllables, and all described constructs are only allowed because the package explicitly implements them.

§ 1 Accents:

       Example: \so{na\"\i␣ve}

Accents can be used naturally. Support for the following accents is built-in: \`, \', \^, \", \~, \=, \., \u, \v, \H, \t, \c, \d, \b, and \r. Additionally, if the german package [8] is loaded you can also use the " accent command and write \so{na"ive}. See section 5.1 for how to add further accents.

4See § 25 for some additional information about the latter mode.

5The look of highlighting is nowhere demonstrated in this documentation, because it requires a Postscript aware output driver and would come out as ugly black bars on other devices, looking very much like censoring bars. Think of it as the effect of one of those coloring text markers.

6Note that you can also use L^{A}TEX’s color package with Plain TEX. See 6.1 for details.

§ 2 Quotes:

Example: \so{``quotes''}

The soul package recognizes the quotes ligatures ``, '' and ,,. The Spanish ligatures !` and ?` are not recognized and have, thus, to be written enclosed in braces like in \caps{{!`}Hola!}.

§ 3 Mathematics:

Example: \so{foo$x^3$bar}

Mathematic formulas are allowed, as long as they are surrounded by $. Note that the L^{A}TEX equivalent \(...\) does not work.

§ 4 Hyphens and dashes:

Example: \so{re-sent}

Explicit hyphens as well as en-dashes (--), em-dashes (---) and the \slash command work as usual.

§ 5 Newlines:

Example: \so{new\\line}

The \\ command fills the current line with white space and starts a new line. Spaces or linebreaks afterwards are ignored. Unlike the original L^{A}TEX command soul’s version does not handle optional parameters like

in \\*[1ex].

§ 6 Breaking lines:

Example: \so{foo\linebreak␣bar}

The \linebreak command breaks the line without filling it with white space at the end. soul’s version does not handle optional parameters like in \linebreak[1]. \break can be used as a synonym.

§ 7 Unbreakable spaces:

Example: \so{don't~break}

The ~ command sets an unbreakable space.

§ 8 Grouping:

Example: \so{Virchow{sche}}

A pair of braces can be used to let a group of characters be seen as one entity, so that soul does for instance not space it out. The contents must, however, not contain potential hyphenation points. (See § 9)

§ 9 Protecting:

Example: \so{foo␣\mbox{little}␣bar}

An \mbox also lets soul see the contents as one item, but these may even contain hyphenation points. \hbox can be used as a synonym.

§ 10 Omitting:

Example: \so{\soulomit{foo}}

The contents of \soulomit bypass the soul core and are typeset as is, without being letterspaced or underlined. Hyphenation points are allowed within the argument. The current font remains active, but can be overridden with \normalfont etc.

§ 11 Font switching commands:

Example: \so{foo␣\texttt{bar}}

All standard TEX and L^{A}TEX font switching commands are allowed, as well as the yfonts package [9] font commands like \textfrak etc. Further commands have to be registered using the \soulregister command (see section 5.2).

§ 12 Breaking up ligatures:

Example: \ul{Auf{}lage}

Use {} or \null to break up ligatures like ‘fl’ in \ul, \st and \hl arguments. This doesn’t make sense for \so and \caps, though, because they break up every unprotected (ungrouped/unboxed) ligature, anyway, and would then just add undesirable extra space around the additional item.

2.2 … others don’t

Although the new soul is much more robust and forgiving than versions prior to 2.0, there are still some things that are not allowed in arguments. This is due to the complex engine, which has to read and inspect every character before it can hand it over to TEX’s paragraph builder.

§ 20 Grouping hyphenatable material:

Example: \so{foo␣{little}␣bar}

Grouped characters must not contain hyphenation points. Instead of

\so{foo {little}} write \so{foo \mbox{little}}. You get a `Re-

construction failed' error and a black square like in the DVI file where you violated this rule.

§ 21 Discretionary hyphens:

Example: \so{Zu\discretionary{k-}{}{c}ker}

The argument must not contain discretionary hyphens. Thus, you have to handle cases like the German word Zu\discretionary{k-}{}{c}ker by yourself.

§ 22 Nested soul commands:

Example: \ul{foo␣\so{bar}␣baz}

soul commands must not be nested. If you really need such, put the inner stuff in a box and use this box. It will, of course, not get broken then.

    \newbox\anyboxname
    \sbox\anyboxname{ \so{the worst} }
    \ul{This is by far{\usebox\anyboxname}example!}

yields:

This is by far t h e w o r s t example!

§ 23 Leaking font switches:

Example: \def\foo{\bf␣bar}␣\so{\foo␣baz}

A hidden font switching command that leaks into its neighborship causes a `Reconstruction failed' error. To avoid this either register the ‘container’ (\soulregister{\foo}{0}), or limit its scope as in \def\foo{{\bf bar}}. Note that both solutions yield slightly different results.

§ 24 Material that needs expansion:

       Example: \so{\romannumeral\year}

In this example \so would try to put space between \romannumeral and \year, which can, of course, not work. You have to expand the argument before you feed it to soul, or even better: Wrap the material up in a com-

       mand sequence and let soul expand it: \def\x{\romannumeral\year}

\so\x. soul tries hard to expand enough, yet not too much.

§ 25 Unexpandable material in command sequences:

       Example: \def\foo{\bar}␣\so\foo

Some macros might not be expandable in an \edef definition^{7} and have to be protected with \noexpand in front. This is automatically done for

       the following tokens: ~, \,, \TeX, \LaTeX, \S, \slash, \textregistered,

\textcircled, and \copyright, as well as for all registered fonts and accents. Instead of putting \noexpand manually in front of such com-

       mands, as in \def\foo{foo {\noexpand\bar} bar} \so\foo, you can

also register them as special (see section 5.2).

§ 26 Other weird stuff:

       Example: \so{foo␣\verb|\bar|␣baz}

soul arguments must not contain L^{A}TEX environments, command definitions, and fancy stuff like \vadjust. soul’s \footnote command replacement does not support optional arguments. As long as you are writing simple, ordinary ‘horizontal’ material, you are on the safe side.

2.3 Troubleshooting

Unfortunately, there’s just one helpful error message provided by the soul package, that actually describes the underlying problem. All other messages are generated directly by TEX and show the low-level commands that TEX wasn’t happy with. They’ll hardly point you to the violated rule as described in the paragraphs above. If you get such a mysterious error message for a line that contains a soul statement, then comment that statement out and see if the message still appears. `Incomplete \ifcat' is such a non-obvious message. If the message doesn’t appear now, then check the argument for violations of the rules as listed in §§ 20–26.

2.3.1 `Reconstruction failed'

This message appears, if § 20 or § 23 was violated. It is caused by the fact that the reconstruction pass couldn’t collect tokens with an overall width of the syllable that was measured by the analyzer. This does either occur when you grouped hyphenatable text or used an unregistered command that influences the syllable width. Font switching commands belong to the latter group. See the above cited sections for how to fix these problems.

2.3.2 Missing characters

If you have redefined the internal font as described in section 5.3, you may notice that some characters are omitted without any error message being shown.

7Try \edef\x{\copyright}. Yet \copyright works in soul arguments, because it is explicitly taken care of by the package

page

\so{letterspacing} 8 l e t t e r s p a c i n g
\caps{CAPITALS, Small Capitals} 10 CAPITALS, Small Capitals
 \ul{underlining} 11 underlining
\st{striking out} 11 striking out
\hl{highlighting} 11 highlighting

\soulaccent{\cs} 14 add accent \cs to accent list

\soulregister{\cs}{0} 15 register command \cs

\sloppyword{text} 19 typeset text with stretchable spaces

\sodef\cs{1em}{2em}{3em} 8 define new spacing command \cs

\resetso 8 reset \so dimensions

\capsdef{////}{1em}{2em}{3em}^{∗} 10 define (default) \caps data entry
  \capssave{name}^{∗} 10 save \caps database under name name
\capsselect{name}^{∗} 10 restore \caps database of name name

\capsreset^{∗} 10 clear caps database

 \setul{1ex}{2ex} 12 set \ul dimensions

\resetul 12 reset \ul dimensions
\setuldepth{y} 12 set underline depth to depth of an y
\setuloverlap{1pt} 13 set underline overlap width
\setulcolor{red} 12 set underline color
\setstcolor{green} 13 set overstriking color
\sethlcolor{blue} 13 set highlighting color

Table 1: List of all available commands. The number points to the page where the command is described. Those marked with a little asterisk are only available when the package is used together with L^{A}TEX, because they rely on the New Font Selection Scheme (NFSS) used in L^{A}TEX.

This happens if you have chosen, let’s say, a font with only 128 characters like the cmtt10 font, but are using characters that aren’t represented in this font, e. g. characters with codes greater than 127.

3 L e t t e r s p a c i n g

3.1 How it works

\so The base macro for letterspacing is called \so. It typesets the given argument with inter-letter space between every two characters, inner space between words and outer space before and after the spaced out text. If we let “·” stand for interletter space, “∗” for inner spaces and “•” for outer spaces, then the input on the left side of the following table will yield the schematic output on the right side:

1.  XX\so{aaa␣bbb␣ccc}YY  XXa·a·a∗b·b·b∗c·c·cYY
2.  XX␣\so{aaa␣bbb␣ccc}␣YY  XX•a·a·a∗b·b·b∗c·c·c•YY
3.  XX␣{\so{aaa␣bbb␣ccc}}␣YY  XX•a·a·a∗b·b·b∗c·c·c•YY

     4. XX␣\null{\so{aaa␣bbb␣ccc}}{}␣YY XX␣a·a·a∗b·b·b∗c·c·c␣YY

Case 1 shows how letterspacing macros (\so and \caps) behave if they aren’t following or followed by a space: they omit outer space around the soul statement. Case 2 is what you’ll mostly need—letterspaced text amidst running text. Following and leading space get replaced by outer space. It doesn’t matter if there are opening braces before or closing braces afterwards. soul can see through both of them (case 3). Note that leading space has to be at least 5sp wide to be recognized as space, because L^{A}TEX uses tiny spaces generated by \hskip1sp as marker. Case 4 shows how to enforce normal spaces instead of outer spaces: Preceding space can be hidden by \kern0pt or \null or any character. Following space can also be hidden by any token, but note that a typical macro name like \relax or \null would also hide the space thereafter.

The values are predefined for typesetting facsimiles mainly with Fraktur fonts. You can define your own spacing macros or overwrite the original \so meaning

\sodef using the macro \sodef:

\sodef⟨cmd⟩{⟨font⟩}{⟨inter-letter space⟩}{⟨inner space⟩}{⟨outer space⟩}

The space dimensions, all of which are mandatory, should be defined in terms of em letting them grow and shrink with the respective fonts.

    \sodef\an{}{.4em}{1em plus1em}{2em plus.1em minus.1em}

\resetso After that you can type ‘\an{example}’ to get ‘e x a m p l e’. The \resetso command resets \so to the default values.

3.2 Some examples

Ordinary text.

                             \so{electrical␣industry}

e l e c t r i c a l i n d u s t r y e l e c - t r i - c a l i n - d u s - t r y Use \- to mark hyphenation points.

                             \so{man\-u\-script}

m a n u s c r i p t

m a n - u - s c r i p t

Accents are recognized.

                             \so{le␣th\'e\^atre}

l e t h é â t r e

l e t h é â t r e

\mbox and \hbox protect

material that contains hyphenation points. The contents are treated as one, unbreakable entity.

                             \so{just␣an␣\mbox{example}}

j u s t a n example

j u s t a n example

Punctuation marks are spaced out, if they are put into the group.

                             \so{inside.}␣\&␣\so{outside}.

i n s i d e . & o u t s i d e.

i n - s i d e . & o u t - s i d e.

Space-out skips may be removed by typing \<. It’s, however, desirable to put the quotation marks out of the argument.

                             \so{``\<Pennsylvania\<''}

“P e n n s y l v a n i a”

“P e n n - s y l - v a - n i a”

Numbers should never be spaced out.

                             \so{1\<3␣December␣{1995}}

13 D e c e m b e r 1995

13 D e - c e m - b e r 1995

Explicit hyphens like -, -- and --- are recognized. \slash outputs a slash and enables TEX to break the line afterwards.

                             \so{input\slash␣output}

i n p u t / o u t p u t

i n - p u t / o u t - p u t

To keep TEX from breaking lines between the hyphen and ‘jet’ you have to protect the hyphen. This is no soul restriction but normal TEX behavior.

                             \so{\dots␣and␣\mbox{-}jet}

… a n d - j e t

… a n d - j e t

The ~ command inhibits line breaks.

                             \so{unbreakable~space}

u n b r e a k a b l e s p a c e u n - b r e a k - a b l e s p a c e

\\ works as usual. Additional arguments like * or vertical space are not accepted, though.

                             \so{broken\\line}

b r o k e n l i n e b r o - k e n l i n e \break breaks the line without filling it with white space.

                             \so{pretty␣awful\break␣test}

p r e t t y a w f u l t e s t

p r e t t y a w - f u l t e s t

3.3 Typesetting capitals-and-small-capitals fonts

\caps There is a special letterspacing command called \caps, which differs from \so in that it switches to caps-and-small-caps font shape, defines only slight spacing and is able to select spacing value sets from a database. This is a requirement for high-quality typesetting [10]. The following lines show the effect of \caps in comparison with the normal textfont and with small-capitals shape:

\normalfont DONAUDAMPFSCHIFFAHRTSGESELLSCHAFT
\scshape DONAUDAMPFSCHIFFAHRTSGESELLSCHAFT
\caps DONAUDAMPFSCHIFFAHRTSGESELLSCHAFT

The \caps font database is by default empty, i. e., it contains just a single default entry, which yields the result as shown in the example above. New font entries

\capsdef may be added on top of this list using the \capsdef command, which takes five arguments: The first argument describes the font with encoding, family, series, shape, and size,^{8} each optionally (e. g. OT1/cmr/m/n/10 for this very font, or only /ppl///12 for all palatino fonts at size 12 pt). The size entry may also contain a size range (5-10), where zero is assumed for an omitted lower boundary (-10) and a very, very big number for an omitted upper boundary (5-). The upper boundary is not included in the range, so, in the example below, all fonts with sizes greater or equal 5 pt and smaller than 15 pt are accepted (5 pt ≤ size < 15 pt). The second argument may contain font switching commands such as \scshape, it may as well be empty or contain debugging commands (e. g. \message{*}). The remaining three, mandatory arguments are the spaces as described in section 3.1.

    \capsdef{T1/ppl/m/n/5-15}{\scshape}{.16em}{.4em}{.2em}

The \caps command goes through the data list from top to bottom and picks up the first matching set, so the order of definition is essential. The last added entry is examined first, while the pre-defined default entry will be examined last and will match any font, if no entry was taken before.

To override the default values, just define a new default entry using the identifier {////}. This entry should be defined first, because no entry after it can be reached.

\capsreset The \caps database can be cleared with the \capsreset command and will \capssave only contain the default entry thereafter. The \capssave command saves the

\capsselect whole current database under the given name. \capsselect restores such a database. This allows to predefine different groups of \caps data sets:

    \capsreset
    \capsdef{/cmss///12}{}{12pt}{23pt}{34pt}
    \capsdef{/cmss///}{}{1em}{2em}{3em}
    ...
    \capssave{wide}

8as defined by the NFSS, the “New Font Selection Scheme”

    %---------------------------------------
    \capsreset
    \capsdef{/cmss///}{}{.1em}{.2em}{.3em}
    ...
    \capssave{narrow}
    %---------------------------------------
    {\capsselect{wide}
    \title{\caps{Yet Another Silly Example}}
    }

See the ‘example.cfg’ file for a detailed example. If you have defined a bunch of sets for different fonts and sizes, you may lose control over what fonts are used

capsdefault by the package. With the package option capsdefault selected, \caps prints its argument underlined, if no set was specified for a particular font and the default set had to be used.

3.4 Typesetting Fraktur

Black letter fonts^{9} deserve some additional considerations. As stated in section 1, the ligatures ch, ck, sz (\ss), and tz have to remain unbroken in spaced out Fraktur text. This may look strange at first glance, but you’ll get used to it:

    \textfrak{\so{S{ch}u{tz}vorri{ch}tung}}

You already know that grouping keeps the soul mechanism from separating such ligatures. This is quite important for s:, a*, and "a. As hyphenation is stronger than grouping, especially the sz may cause an error, if hyphenation happens to occur between the letters s and z. (TEX hyphenates the German word auszer wrongly like aus-zer instead of like au-szer, because the German hyphenation patterns do, for good reason, not see sz as ‘\ss’.) In such cases you can protect tokens with the sequence e. g. \mbox{sz} or a properly defined command. The \ss command, which is defined by the yfonts package, and similar commands will suffice as well.

3.5 Dirty tricks

Narrow columns are hard to set, because they don’t allow much spacing flexibility, hence long words often cause overfull boxes. A macro could use \so to insert stretchability between the single characters. Table 2 shows some text typeset with such a macro at the left side and under plain conditions at the right side, both with a width of 6 pc.

4 Underlining

The underlining macros are my answer to Prof. Knuth’s exercise 18.26 from his \ul TEXbook [5]. :-) Most of what is said about the macro \ul is also true of the \st striking out macro \st and the highlighting macro \hl, both of which are in fact \hl derived from the former.

9See the great black letter fonts, which Yannis Haralambous kindly provided, and the oldgerm and yfonts package [9] as their L^{A}TEX interfaces.

Some magazines and newspapers prefer this kind of spacing because it reduces hyphenation problems to a minimum. Unfortunately, such paragraphs aren’t especially beautiful.

Some magazines and newspapers prefer this kind of spacing b ecause it reduces h y p h e n a t i o n p r o b l e m s t o a minimum. Unf o r t u n a t e l y, such paragraphs aren’t especially beautiful.

Some magazines and newspapers prefer this kind of spacing because it reduces hyphenation problems to a minimum. Unfortunately, such paragraphs aren’t especially beautiful.

Table 2: Ragged-right, magazine style (using soul), and block-aligned in comparison. But, frankly, none of them is really acceptable. (Don’t do this at home, children!)

4.1 Settings

4.1.1 Underline depth and thickness

The predefined underline depth and thickness work well with most fonts. They \setul can be changed using the macro \setul.

\setul{⟨underline depth⟩}{⟨underline thickness⟩}

Either dimension can be omitted, in which case there has to be an empty pair of braces. Both values should be defined in terms of ex, letting them grow and

\resetul shrink with the respective fonts. The \resetul command restores the standard

values.

\setuldepth Another way to set the underline depth is to use the macro \setuldepth. It sets the depth such that the underline’s upper edge lies 1 pt beneath the given argument’s deepest depth. If the argument is empty, all letters—i. e. all characters whose \catcode currently equals 11—are taken. Examples:

    \setuldepth{ygp}
    \setuldepth\strut
    \setuldepth{}

4.1.2 Line color

The underlines are by default black. The color can be changed by using the

\setulcolor \setulcolor command. It takes one argument that can be any of the color specifiers as described in the color package. This package has to be loaded explicitly.

    \documentclass{article}
    \usepackage{color,soul}
    \definecolor{darkblue}{rgb}{0,0,0.5}
    \setulcolor{darkblue}

    \begin{document}
    ...
    \ul{Cave: remove all the underlines!}
    ...
    \end{document}

\setstcolor The colors for overstriking lines and highlighting are likewise set with \setstcolor \sethlcolor (default: black) and \sethlcolor (default: yellow). If the color package wasn’t

loaded, underlining and overstriking color are black, while highlighting is replaced by underlining.

4.1.3 The dvips problem

Underlining, striking out and highlighting build up their lines with many short line segments. If you used the ‘dvips’ program with default settings, you would get little gaps on some places, because the maxdrift parameter allows the single objects to drift this many pixels from their real positions.

There are two ways to avoid the problem, where the soul package chooses the second by default:

1. Set the maxdrift value to zero, e. g.: dvips -e 0 file.dvi. This is probably not a good idea, since the letters may then no longer be spaced equally on low resolution printers.

2. Let the lines stick out by a certain amount on each side so that they overlap.

\setuloverlap This overlap amount can be set using the \setuloverlap command. It is set to 0.25 pt by default. \setuloverlap{0pt} turns overlapping off.

4.2 Some examples

Ordinary text.

                             \ul{electrical␣industry}

electrical industry electrical industry

Use \- to mark hyphenation points.

                             \ul{man\-u\-script}

manuscript

man- u- script

Accents are recognized.

                             \ul{le␣th\'e\^atre}

le théâtre

le théâtre

\mbox and \hbox protect

material that contains hyphenation points. The contents are treated as one, unbreakable entity.

                             \ul{just␣an␣\mbox{example}}

just an example just an example Explicit hyphens like -, -- and --- are recognized. \slash outputs a slash and enables TEX to break the line afterwards.

                             \ul{input\slash␣output}

input/output

in- put/ out- put

To keep TEX from breaking lines between the hyphen and ‘jet’ you have to protect the hyphen. This is no soul restriction but normal TEX behavior.

                             \ul{\dots␣and␣\mbox{-}jet}

… and -jet

… and -jet

The ~ command inhibits line breaks.

                             \ul{unbreakable~space}

unbreakable space

un- break- able space

\\ works as usual. Additional arguments like * or vertical space are not accepted, though.

                             \ul{broken\\line}

broken line

bro- ken line

\break breaks the line without filling it with white space.

                             \ul{pretty␣awful\break␣test}

pretty awful test

pretty aw- ful test

Italic correction needs to be set manually.

                             \ul{foo␣\emph{bar\/}␣baz}

foo bar baz

foo bar baz

5 Customization

5.1 Adding accents

The soul scanner generally sees every input token separately. It has to be taught that some tokens belong together. For accents this is done by registering them via

\soulaccent the \soulaccent macro.

\soulaccent{⟨accent command⟩}

The standard accents, however, are already pre-registered: \`, \', \^, \", \~, \=, \., \u, \v, \H, \t, \c, \d, \b, and \r. If used together with the german package, soul automatically adds the " command. Let’s assume you have defined \% to put some weird accent on the next character. Simply put the following line into your soul.cfg file (see section 5.4):

    \soulaccent{\%}

Note that active characters like the " command have already to be \active when they are stored or they won’t be recognized later. This can be done temporarily,

as in {\catcode\`"\active\soulaccent{"}}.

5.2 Adding font commands

To convince soul not to feed font switching (or other) commands to the analyzer, but rather to execute them immediately, they have to be registered, too. The

\soulregister \soulregister macro takes the name of a command name and either 0 or 1 for the number of arguments:

\soulregister{⟨command name⟩}{⟨number of arguments⟩}

If \bf and \emph weren’t already registered, you would write the following into your soul.cfg configuration file:

    \soulregister{\bf}{0}        % {\bf foo}
    \soulregister{\emph}{1}      % \emph{bar}

All standard TEX and L^{A}TEX font commands, as well as the yfonts commands are already pre-registered:

    \em, \rm, \bf, \it, \tt, \sc, \sl, \sf, \emph, \textrm,
    \textsf, \texttt, \textmd, \textbf, \textup, \textsl,
    \textit, \textsc, \textnormal, \rmfamily, \sffamily,
    \ttfamily, \mdseries, \upshape, \slshape, \itshape,
    \scshape, \normalfont, \tiny, \scriptsize, \footnotesize,
    \small, \normalsize, \large, \Large, \LARGE, \huge, \Huge,
    \MakeUppercase, \textsuperscript, \footnote,
    \textfrak, \textswab, \textgoth, \frakfamily,
    \swabfamily, \gothfamily

You can also register other commands as fonts, so the analyzer won’t see them. This may be necessary for some macros that soul refuses to typeset correctly. But note, that \so and \caps won’t put their letter-skips around then.

5.3 Changing the internal font

The soul package uses the ectt1000 font while it analyzes the syllables. This font is used, because it has 256 mono-spaced characters without any kerning. It belongs to Jörg Knappen’s EC-fonts, which should be part of every modern TEX installation. If TEX reports “I can't find file `ectt1000'” you don’t seem to have this font installed. It is recommended that you install at least the file ectt1000.tfm which has less than 1.4 kB. Alternatively, you can let the soul package use the cmtt10 font that is part of any installation, or some other monospaced font:

    \font\SOUL@tt=cmtt10

Note, however, that soul does only handle characters, for which the internal font has a character with the same character code. As cmtt10 contains only characters with codes 0 to 127, you can’t typeset characters with codes 128 to 255. These 8-bit character codes are used by many fonts with non-ascii glyphs. So the cmtt10 font will, for example, not work for T2A encoded cyrillic characters.

5.4 The configuration file

If you want to change the predefined settings or add new features, then create a file named ‘soul.cfg’ and put it in a directory, where TEX can find it. This configuration file will then be loaded at the end of the soul.sty file, so you may redefine any settings or commands therein, select package options and even introduce new ones. But if you intend to give your documents to others, don’t forget to give them the required configuration files, too! That’s how such a file could look like:

    % define macros for logical markup
    \sodef\person{\scshape}{0.125em}{0.4583em}{0.5833em}
    \sodef\SOUL@@@versal{\upshape}{0.125em}{0.4583em}{0.5833em}
    \DeclareRobustCommand*\versal[1]{%
        \MakeUppercase{\SOUL@@@versal{#1}}%
    }
    % load the color package and set
    % a different highlighting color
    \RequirePackage{color}
    \definecolor{lightblue}{rgb}{.90,.95,1}
    \sethlcolor{lightblue}
    \endinput

You can safely use the \SOUL@@@ namespace for internal macros—it won’t be used by the soul package in the future.

6 Miscellaneous

6.1 Using soul with other flavors of TEX

This documentation describes how to use soul together with L^{A}TEX 2ε, for which it is optimized. It works, however, with all other flavors of TEX, too. There are just some minor restrictions for Non-L^{A}TEX use:

The \caps command doesn’t use a database, it is only a dumb definition with fixed values. It switches to \capsfont, which—unless defined explicitly like in the following example—won’t really change the used font at all. The commands

\capsreset and \capssave do nothing.
    \font\capsfont=cmcsc10
    \caps{Tschichold}

None of the commands are made ‘robust’, so they have to be explicitly protected in fragile environments like in \write statements. To make use of colored underlines or highlighting you have to use the color package wrapper from CTAN^{10}, instead of the color package directly:

    \input color
    \input soul.sty
    \hl{highlighted}
    \bye

\capsdefault The capsdefault package option is mapped to a simple command \capsdefault.

10CTAN:/macros/plain/graphics/{miniltx.tex,color.tex}

6.2 Using soul commands for logical markup

It’s generally a bad idea to use font style commands like \textsc in running text. There should always be some reasoning behind changing the style, such as “names of persons shall be typeset in a caps-and-small-caps font”. So you declare in your text just that some words are the name of a person, while you define in the preamble or, even better, in a separate style file how to deal with persons:

    \newcommand*\person{\textsc}
    ...
    ``I think it's a beautiful day to go to the zoo and feed
    the ducks. To the lions.'' --~\person{Brian Kantor}

It’s quite simple to use soul commands that way:

    \newcommand\comment*{\ul}             % or \let\comment=\ul
    \sodef\person{\scshape}{0.125em}{0.4583em}{0.5833em}

Letterspacing commands like \so and \caps have to check whether they are followed by white space, in which case they replace that space by outer space. Note that soul does look through closing braces. Hence you can conveniently bury a soul command within another macro like in the following example. Use any other token to hide following space if necessary, for example the \null macro.

    \DeclareRobustCommand*\versal[1]{%
        \MakeUppercase{\SOUL@@@versal{#1}}%
    }
    \sodef\SOUL@@@versal{\upshape}{0.125em}{0.4583em}{0.5833em}

But what if the soul command is for some reason not the last one in that macro definition and thus cannot look ahead at the following token?

    \newcommand*\somsg[1]{\so{#1}\message{#1}}
    ...
    foo \somsg{bar} baz       % wrong spacing after `bar'!

In this case you won’t get the following space replaced by outer space because when soul tries to look ahead, it only sees the token \message and consequently decides that there is no space to replace. You can get around this by explicitly calling the space scanner again.

    \newcommand*\somsg[1]{{%
        \so{#1}%
        \message{bar}%
        \let\\\SOUL@socheck
        \\%
    }}

However, \SOUL@socheck can’t be used directly, because it would discard any normal space. \\ doesn’t have this problem. The additional pair of braces avoids that its definition leaks out of this macro. In the example above you could, of course, simply have put \message in front, so you hadn’t needed to use the scanner

macro \SOUL@socheck at all.

Many packages do already offer logical markup commands that default to some standard L^{A}TEX font commands or to \relax. One example is the jurabib package [1], which makes the use of soul a challenge. This package implements lots of formatting macros. Let’s have a look at one of them, \jbauthorfont, which is used to typeset author names in citations. The attempt to simply define \let\jbauthorfont\caps fails, because the macro isn’t directly applied to the author name as in \jbauthorfont{Don Knuth}, but to another command sequence:

\jbauthorfont{\jb@@author}. Not even \jb@@author contains the name, but

instead further commands that at last yield the requested name. That’s why we have to expand the contents first. This is quite tricky, because we must not expand too much, either. Fortunately, we can offer the contents wrapped up in yet another macro, so that soul knows that it has to use its own macro expansion mechanism:

    \renewcommand*\jbauthorfont[1]{{%
        \def\x{#1}%
        \caps\x
    }}

Some additional kerning after \caps\x wouldn’t hurt, because the look-ahead scanner is blinded by further commands that follow in the jurabib package. Now we run into the next problem: cited names may contain commands that must not get expanded. We have to register them as special command:

    \soulregister\jbbtasep{0}
    ...

But such registered commands bypass soul’s kernel and we don’t get the correct spacing before and afterwards. So we end up redefining \jbbtasep, whereby you should, of course, use variables instead of numbers:

    \renewcommand*\jbbtasep{%
        \kern.06em
        \slash
        \hskip.06em
        \allowbreak
    }

Another problem arises: bibliography entries that must not get teared apart are supposed to be enclosed in additional braces. This, however, won’t work with soul because of § 20. A simple trick will get you around that problem: define a dummy command that only outputs its argument, and register that command:

    \newcommand*\together[1]{#1}
    \soulregister\together{1}
Now you can write “Author = {\together{Don Knuth}}” and jurabib won’t

dare to reorder the parts of the name. And what if some name shouldn’t get letterspaced at all? Overriding a conventional font style like \textbf that was globally set is trivial, you just have to specify the style that you prefer in that very bibliography entry. In our example, if we wanted to keep soul from letterspacing a particular entry, although they are all formatted by our \jbauthorfont and hence fed to \caps, we’d use the following construction:

    Author = {\soulomit{\normalfont\huge Donald E. Knuth}}

The jurabib package is probably one of the more demanding packages to collaborate with soul. Everything else can just become easier.

6.3 Typesetting long words in narrow columns

Narrow columns are best set flushleft, because not even the best hyphenation algorithm can guarantee acceptable line breaks without overly stretched spaces. However, in some rare cases one may be forced to typeset block aligned. When typesetting in languages like German, where there are really long words, the

\sloppyword \sloppyword macro might help a little bit. It adds enough stretchability between the single characters to make the hyphenation algorithm happy, but is still not as ugly as the example in section 3.5 demonstrates. In the following example the left

column was typeset as “Die \sloppyword{Donau...novelle} wird ...”:

Die Donaudampfschifffahrtsgesellschaftskapitänswitwenpensionsgesetznovelle wird mit sofortiger Wirkung außer Kraft gesetzt.

Die Donaudampfschifffahrtsgesellschaftskapitänswitwenpensionsgesetznovelle wird mit sofortiger Wirkung außer Kraft gesetzt.

6.4 Using soul commands in section headings

Letterspacing was often used for section titles in the past, mostly centered and with a closing period. The following example shows how to achieve this using the titlesec package [2]:

    \newcommand*\periodafter[2]{#1{#2}.}
    \titleformat{\section}[block]
        {\normalfont\centering}
        {\thesection.}
        {.66em}
        {\periodafter\so}
    ...
    \section{Von den Maassen und Maassst\"aben}

This yields the following output:

1. V o n d e n M a a s s e n u n d M a a s s s t ä b e n.

The \periodafter macro adds a period to the title, but not to the entry in the table of contents. It takes the name of a command as argument, that shall be applied to the title, for example \so. Here’s a more complicated and complete example:

    \documentclass{article}
    \usepackage[latin1]{inputenc}
    \usepackage[T1]{fontenc}
    \usepackage{german,soul}
    \usepackage[indentfirst]{titlesec}

\newcommand*\sectitle[1]{%
    \MakeUppercase{\so{#1}.}\\[.66ex]
    \rule{13mm}{.4pt}}
\newcommand*\periodafter[2]{#1{#2.}}
\titleformat{\section}[display]
    {\normalfont\centering}
    {\S. \thesection.}
    {2ex}
    {\sectitle}
\titleformat{\subsection}[block]
    {\normalfont\centering\bfseries}
    {\thesection.}
    {.66em}
    {\periodafter\relax}
\begin{document}
\section{Von den Maassen und Maassst\"aben}
\subsection{Das L\"angenmaass im Allgemeinen}
Um L\"angen genau messen und vergleichen zu k\"onnen,
bedarf es einer gewissen, bestimmten Einheit, mit der
man untersucht, wie oft sie selbst, oder ihre Theile,
in der zu bestimmenden L\"ange enthalten sind.
...
\end{document}

This example gives you roughly the following output, which is a facsimile from [6].

§. 1.

V O N D E N M A A S S E N U N D M A A S S S T Ä B E N.

1. Das Längenmaass im Allgemeinen.

Um Längen genau messen und vergleichen zu können, bedarf es einer gewissen, bestimmten Einheit, mit der man untersucht, wie oft sie selbst, oder ihre Theile, in der zu bestimmenden Länge enthalten sind.

Note that the definition of \periodafter decides if the closing period shall be spaced out with the title (1), or follow without space (2):

1.  \newcommand*\periodafter[2]{#1{#2.}}
2.  \newcommand*\periodafter[2]{#1{#2}.}

If you need to underline section titles, you can easily do it with the help of the titlesec package. The following example underlines the section title, but not the section number:

    \titleformat{\section}
        {\LARGE\titlefont}
        {\thesection}
        {.66em}
        {\ul}

The \titlefont command is provided by the “KOMA script” package. You can write \normalfont\sffamily\bfseries instead. The following example does additionally underline the section number:

    \titleformat{\section}
        {\LARGE\titlefont}
        {\ul{\thesection{\kern.66em}}}
        {0pt}
        {\ul}

7 How the package works

7.1 The kernel

L e t t e r s p a c i n g , underlining, striking out and highlighting use the same kernel. It lets a word scanner run over the given argument, which inspects every token. If a token is a command registered via \soulregister, it is executed immediately. Other tokens are only counted and trigger some action when a certain number is reached (quotes and dashes). Three subsequent ‘-’, for example, trigger \SOUL@everyexhyphen{---}. A third group leads to special actions, like \mbox that starts reading-in a whole group to protect its contents and let them be seen as one entity. All other tokens, mostly characters and digits, are collected in a word register, which is passed to the analyzer, whenever a whole word was read in.

The analyzer typesets the word in a 1 sp (= \frac{1}{65536} pt) wide \vbox, hence encouraging TEX to break lines at every possible hyphenation point. It uses the mono-spaced \SOUL@tt font (ectt1000), so as to avoid any inter-character kerning. Now the \vbox is decomposed splitting off \hbox after \hbox from the bottom. All boxes, each of which contains one syllable, are pushed onto a stack, which is provided by TEX’s grouping mechanism. When returning from the recursion, box after box is fetched from the stack, its width measured and fed to the “reconstructor”.

This reconstruction macro (\SOUL@dosyllable) starts to read tokens from the just analyzed word until the given syllable width is obtained. This is repeated for each syllable. Every time the engine reaches a relevant state, the corresponding driver macro is executed and, if necessary, provided with some data. There is a macro that is executed for each token, one for each syllable, one for each space etc.

The engine itself doesn’t know how to letterspace or to underline. It just tells the selected driver about the structure of the given argument. There’s a default driver (\SOUL@setup) that does only set the interface macros to a reasonable default state, but doesn’t really do anything. Further drivers can safely inherit these settings and only need to redefine what they want to change.

7.2 The interface

7.2.1 The registers

The package offers eight interface macros that can be used to define the required actions. Some of the macros receive data as macro parameter or in special token or dimen registers. Here is a list of all available registers:

\SOUL@token This token register contains the current token. It has to be used as \the\SOUL@token. The macro \SOUL@gettoken reads the next token into \SOUL@token and can be used in any interface macro. If you don’t want to lose the old meaning, you have to save it explicitly. \SOUL@puttoken pushes the token back into the queue, without changing \SOUL@token. You can only put one token back, otherwise you get an error message.

\SOUL@lasttoken This token register contains the last token.

\SOUL@syllable This token register contains all tokens that were already collected for the current syllable. When used in \SOUL@everysyllable, it contains the whole syllable.

\SOUL@charkern This dimen register contains the kerning value between the current and the next character. Since most character pairs don’t require a kerning value to be applied and the output in the logfile shouldn’t be cluttered with \kern0pt it is recommended to

write \SOUL@setkern\SOUL@charkern, which sets

kerning for non-zero values only.

\SOUL@hyphkern This dimen register contains the kerning value between the current character and the hyphen character or, when used in \SOUL@everyexhyphen, the kerning between the last character and the explicit hyphen.

7.2.2 The interface macros

The following list describes each of the interface macros and which registers it can rely on. The mark between label and description will be used in section 7.2.3 to show when the macros are executed. The addition #1 means that the macro takes one argument.

\SOUL@preamble P executed once at the beginning

\SOUL@postamble E executed once at the end \SOUL@everytoken T executed after scanning a token; It gets that to-

ken in \SOUL@token and has to care for inserting the kerning value \SOUL@charkern between this and the next character. To look at the next character, execute \SOUL@gettoken, which replaces \SOUL@token by the next token. This token has to be put back into the queue using \SOUL@puttoken.

\SOUL@everysyllable S This macro is executed after scanning a whole syllable. It gets the syllable in \SOUL@syllable.

\SOUL@everyhyphen − This macro is executed at every implicit hyphenation point. It is responsible for setting the hyphen and will likely do this in a \discretionary statement. It has to care about the kerning values.

The registers \SOUL@lasttoken, \SOUL@syllable,
\SOUL@charkern and \SOUL@hyphkern contain use-

ful information. Note that \discretionary inserts \exhyphenpenalty if the first part of the discretionary is empty, and \hyphenpenalty else.

\SOUL@everyexhyphen#1 = This macro is executed at every explicit hyphenation point. The hyphen ‘character’ (one of hyphen, en-dash, em-dash or \slash) is passed as parameter #1. A minimal implementation would

be {#1\penalty\exhyphenpenalty}. The kerning

value between the last character and the hyphen is passed in \SOUL@hyphkern, that between the hyphen and the next character in \SOUL@charkern. The last syllable can be found in \SOUL@syllable, the last character in \SOUL@lasttoken.

\SOUL@everyspace#1 ␣ This macro is executed between every two words. It is responsible for setting the space. The engine submits a \penalty setting as parameter #1 that should be put in front of the space. The macro should at least do {#1\space}. Further information can be found in \SOUL@lasttoken and \SOUL@syllable. Note that this macro does not care for the leading and trailing space. This is the

job of \SOUL@preamble and \SOUL@postamble.

7.2.3 Some examples

The above list’s middle column shows a mark that indicates in the following examples, when the respective macros are executed:

P w^{T} o^{T} r^{T} d^{T} ^{S}^{E} \SOUL@everytoken^{T} is executed for every token. \SOUL@everysyllable^{S} is additionally executed for every syllable. You will mostly just want to use either of them.

P o^{T} n^{T} e^{T} ^{S}␣t^{T} w^{T} o^{T} ^{S}^{E} The macro \SOUL@everyspace is executed at every space within the soul argument. It has to take one argument, that can either be empty or contain a penalty, that should be applied to the space.

P e^{T} x^{T} ^{S}^{−}a^{T} m^{T} ^{S}^{−}p^{T} l^{T} e^{T} ^{S}^{E} The macro \SOUL@everyhyphen is executed at every possible implicit hyphenation point.

P b^{T} e^{T} t^{T} a^{T} ^{S}-^{=}t^{T} e^{T} s^{T} t^{T} ^{S}^{E} Explicit hyphens trigger \SOUL@everyexhyphen.

It’s only natural that these examples, too, were automatically typeset by the soul package using a special driver:

    \DeclareRobustCommand*\an{%
        \def\SOUL@preamble{$^{^P}$}%
        \def\SOUL@everyspace##1{##1\texttt{\char`\ }}%
        \def\SOUL@postamble{$^{^E}$}%
        \def\SOUL@everyhyphen{$^{^-}$}%
        \def\SOUL@everyexhyphen##1{##1$^{^=}$}%
        \def\SOUL@everysyllable{$^{^S}$}%
        \def\SOUL@everytoken{\the\SOUL@token$^{^T}$}%
        \def\SOUL@everylowerthan{$^{^L}$}%
        \SOUL@}

7.3 A driver example

Let’s define a soul driver that allows to typeset text with a \cdot at every potential hyphenation point. The name of the macro shall be \sy (for syllables). Since the soul mechanism is highly fragile, we use the L^{A}TEX command \DeclareRobustCommand, so that the \sy macro can be used even in section headings etc. The \SOUL@setup macro sets all interface macros to reasonable default definitions. This could of course be done manually, too. As we won’t make use of

\SOUL@everytoken and \SOUL@postamble and both default to \relax, anyway,

we don’t have to define them here.

    \DeclareRobustCommand*\sy{%
        \SOUL@setup

We only set \lefthyphenmin and \righthyphenmin to zero at the beginning. All changes are restored automatically, so there’s nothing to do at the end.

        \def\SOUL@preamble{\lefthyphenmin=0 \righthyphenmin=0 }%

We only want simple spaces. Note that these are not provided by default! \SOUL@everyspace may get a penalty to be applied to that space, so we set it before.

        \def\SOUL@everyspace##1{##1\space}%
There’s nothing to do for \SOUL@everytoken, we rather let \SOUL@everysyllable

handle a whole syllable at once. This has the advantage, that we don’t have to deal with kerning values, because TEX takes care of that.

        \def\SOUL@everysyllable{\the\SOUL@syllable}

The TEX primitive \discretionary takes three arguments: 1. pre-hyphen material 2. post-hyphen material, and 3. no-hyphenation material.

        \def\SOUL@everyhyphen{%
            \discretionary{%
                \SOUL@setkern\SOUL@hyphkern
                \SOUL@sethyphenchar
            }{}{%
                \hbox{\kern1pt$\cdot$}%
            }%
        }%

Explicit hyphens like dashes and slashes shall be set normally. We just have to care for kerning. The hyphen has to be put in a box, because, as \hyphenchar, it would yield its own, internal \discretionary. We need to set ours instead, though.

        \def\SOUL@everyexhyphen##1{%
            \SOUL@setkern\SOUL@hyphkern
            \hbox{##1}%
            \discretionary{}{}{%
                \SOUL@setkern\SOUL@charkern
            }%
        }

Now that the interface macros are defined, we can start the scanner.

    {\SOUL@ \atop \}}

This lit·tle macro will hard·ly be good e·nough for lin·guists, al·though it us·es TEX’s ex·cel·lent hy·phen·ation al·go·rithm, but it is at least a nice al·ter·na·tive to the \showhyphens com·mand.

Acknowledgements

A big thank you goes to Stefan Ulrich for his tips and bug reports during the development of versions 1.* and for his lessons on high quality typesetting. The \caps mechanism was very much influenced by his suggestions. Thanks to Alexander Shibakov and Frank Mittelbach, who sent me a couple of bug reports and feature requests, and finally encouraged me to (almost) completely rewrite soul. Thorsten Manegold contributed a series of bug reports, helping to fix soul’s macro expander and hence making it work together with the jurabib package. Thanks to Axel Reichert, Anshuman Pandey, and Peter Kreynin for detailed bug reports. Rowland McDonnel gave useful hints for how to improve the documentation, but I’m afraid he will still not be satisfied, and rightfully so. If only documentation writing weren’t that boring. ;-)

References

[1] Berger, Jens: The jurabib package. CTAN-Archive, 2002, v0.52h.

[2] Bezos, Javier: The titlesec and titletoc package. CTAN-Archive, 1999, v2.1.

[3] Carlisle, D. P.: The color package. CTAN-Archive, 1997, v1.0d.

[4] Duden, Volume 1. Die Rechtschreibung. Bibliographisches Institut, Mannheim–Wien–Zürich, 1986, 19th edition.

[5] Knuth, Donald E.: The TEXbook. Addison–Wesley Publishing Company, Reading/Massachusetts, 1989, 16th edition.

[6] Muszynski, Carl and Přihoda, Eduard: Die Terrainlehre in Verbindung mit der Darstellung, Beurtheilung und Beschreibung des Terrains vom militärischen Standpunkte. L. W. Seidel & Sohn, Wien, 1872.

[7] Normalverordnungsblatt für das k. u. k. Heer. Exercier-Reglement für die k. u. k. Cavallerie, I. Theil. Wien, k. k. Hof- und Staatsdruckerei, 1898, 4th edition.

[8] Raichle, Bernd: The german package. CTAN-Archive, 1998, v2.5e.

[9] Schmidt, Walter: Ein Makropaket für die gebrochenen Schriften. CTAN-Archive, 1998, v1.2.

[10] Tschichold, Jan: Ausgewählte Aufsätze über Fragen der Gestalt des Buches und der Typographie. Birkhäuser, Basel, 1987, 2nd edition.

[11] Willberg, Hans Peter and Forssmann, Friedrich:

Lesetypographie. H. Schmidt, Mainz, 1997.

8 The implementation

The package preamble

This piece of code makes sure that the package is only loaded once. While this is guaranteed by L^{A}TEX, we have to do it manually for all other flavors of TEX.

1 \expandafter\ifx\csname SOUL@\endcsname\relax\else
2  \expandafter\endinput
3 \fi

Fake some of the L^{A}TEX commands if we were loaded by another flavor of TEX. This might break some previously loaded packages, though, if e. g. \mbox was already in use. But we don’t care …

4 \ifx\documentclass\SOULundefined
5  \chardef\atcode=\catcode`@
6  \catcode`\@=11
7  \def\DeclareRobustCommand*{\def}
8  \let\newcommand\DeclareRobustCommand
9  \def\DeclareOption#1#2{\expandafter\def\csname#1\endcsname{#2}}
10  \def\PackageError#1#2#3{{%
11  \newlinechar`^^J%
12  \errorcontextlines\z@
13  \edef\\{\errhelp{#3}}\\%

14  \errmessage{Package #1 error: #2}%
15  }}
16  \def\@height{height}
17  \def\@depth{depth}
18  \def\@width{width}
19  \def\@plus{plus}
20  \def\@minus{minus}
21  \font\SOUL@tt=ectt1000
22  \let\@xobeysp\space
23  \let\linebreak\break
24  \let\mbox\hbox

soul tries to be a good L^{A}TEX citizen if used under L^{A}TEX and declares itself properly. Most command sequences in the package are protected by the SOUL@ namespace, all other macros are first defined to be empty. This will give us an error message now if one of those was already used by another package.

25 \else
26  \NeedsTeXFormat{LaTeX2e}
27  \ProvidesPackage{soul}
28  [2003/11/17 v2.4 letterspacing/underlining  (mf)]
29  \newfont\SOUL@tt{ectt1000}
30  \newcommand*\sodef{}
31  \newcommand*\resetso{}
32  \newcommand*\capsdef{}
33  \newcommand*\capsfont{}
34  \newcommand*\setulcolor{}
35  \newcommand*\setuloverlap{}
36  \newcommand*\setul{}
37  \newcommand*\resetul{}
38  \newcommand*\setuldepth{}
39  \newcommand*\setstcolor{}
40  \newcommand*\sethlcolor{}
41  \newcommand*\so{}
42  \newcommand*\ul{}
43  \newcommand*\st{}
44  \newcommand*\hl{}
45  \newcommand*\caps{}
46  \newcommand*\soulaccent{}
47  \newcommand*\soulregister{}
48  \newcommand*\soulfont{}
49  \newcommand*\soulomit{}
50 \fi

Other packages wouldn’t be happy if we reserved piles of \newtoks and \newdimen, so we try to get away with their \...def counterparts where possible. Local registers are always even, while global ones are odd—this is a TEX convention.

51 \newtoks\SOUL@word
52 \newtoks\SOUL@lasttoken
53 \toksdef\SOUL@syllable\z@
54 \newtoks\SOUL@cmds
55 \newtoks\SOUL@buffer
56 \newtoks\SOUL@token
57 \dimendef\SOUL@syllgoal\z@
58 \dimendef\SOUL@syllwidth\tw@

59 \dimendef\SOUL@charkern=4
60 \dimendef\SOUL@hyphkern=6
61 \countdef\SOUL@minus\z@
62 \countdef\SOUL@comma\tw@
63 \countdef\SOUL@apo=4
64 \countdef\SOUL@grave=6
65 \newskip\SOUL@spaceskip
66 \newif\ifSOUL@ignorespaces
\soulomit \SOUL@ignorem \SOUL@ignore \SOUL@stopm \SOUL@stop \SOUL@relaxm \SOUL@lowerthanm \SOUL@hyphenhintm These macros are used as markers. To be able to check for such a marker with

\ifx we have also to create a macro that contains the marker. \SOUL@spc shall contain a normal space with a \catcode of 10.

67 \def\soulomit#1{#1}
68 \def\SOUL@stopM{\SOUL@stop}
69 \let\SOUL@stop\relax
70 \def\SOUL@lowerthan{}
71 \def\SOUL@lowerthanM{\<}
72 \def\SOUL@hyphenhintM{\-}
73 \def\SOUL@n*{\let\SOUL@spc= }\SOUL@n* %

8.1 The kernel

\SOUL@ This macro is the entry to soul. Using it does only make sense after setting up a soul driver. The next token after the soul command will be assigned to \SOUL@@. This can be some text enclosed in braces, or the name of a macro that contains text.

74 \def\SOUL@{%
75  \futurelet\SOUL@@\SOUL@expand

76 }

\SOUL@expand If the first token after the soul command was an opening brace we start scanning. Otherwise, if the first token was a macro name, we expand that macro and call \SOUL@ with its contents again. Unfortunately, we have to exclude some macros therein from expansion.

77 \def\SOUL@expand{%
78  \ifcat\bgroup\noexpand\SOUL@@
79  \let\SOUL@n\SOUL@start
80  \else
81  \bgroup
82  \def\\##1##2{\def##2{\noexpand##2}}%
83  \the\SOUL@cmds
84  \SOUL@buffer={%
85  \\\TeX\\\LaTeX\\\soulomit\\\mbox\\\hbox\\\textregistered
86  \\\slash\\\textcircled\\\copyright\\\S\\\,\\\<\\\>\\~%
87  \\\\%
88  }%
89  \def\\##1{\def##1{\noexpand##1}}%
90  \the\SOUL@buffer
91  \let\protect\noexpand
92  \xdef\SOUL@n##1{\noexpand\SOUL@start{\SOUL@@}}%
93  \egroup
94  \fi

 95  \SOUL@n

96 }

 97 \long\def\SOUL@start#1{{%
 98  \let\<\SOUL@lowerthan
 99  \let\>\empty
100  \def\soulomit{\noexpand\soulomit}%
101  \gdef\SOUL@eventuallyexhyphen##1{}%
102  \let\SOUL@soeventuallyskip\relax
103  \SOUL@spaceskip=\fontdimen\tw@\font\@plus\fontdimen\thr@@\font
104  \@minus\fontdimen4\font
105  \SOUL@ignorespacesfalse
106  \leavevmode
107  \SOUL@preamble
108  \SOUL@lasttoken={}%
109  \SOUL@word={}%
110  \SOUL@minus\z@
111  \SOUL@comma\z@
112  \SOUL@apo\z@
113  \SOUL@grave\z@
114  \SOUL@do{#1}%
115  \SOUL@postamble

116 }}

117 \long\def\SOUL@do#1{%
118  \SOUL@scan#1\SOUL@stop

119 }

8.2 The scanner

\SOUL@scan This is the entry point for the scanner. It calls \SOUL@eval and will in turn be called by \SOUL@eval again for every new token to be scanned.

120 \def\SOUL@scan{%
121  \futurelet\SOUL@@\SOUL@eval

122 }

\SOUL@eval And here it is: the scanner’s heart. It cares for quotes and dashes ligatures and handles all commands that must not be fed to the analyzer.

123 \def\SOUL@eval{%
124  \def\SOUL@n*##1{\SOUL@scan}%
125  \if\noexpand\SOUL@@\SOUL@spc
126  \else
127  \SOUL@ignorespacesfalse
128  \fi
129  \ifnum\SOUL@minus=\thr@@
130  \SOUL@flushminus
131  \else\ifnum\SOUL@comma=\tw@
132  \SOUL@flushcomma
133  \else\ifnum\SOUL@apo=\tw@
134  \SOUL@flushapo
135  \else\ifnum\SOUL@grave=\tw@
136  \SOUL@flushgrave
137  \fi\fi\fi\fi
138  \ifx\SOUL@@-\else\SOUL@flushminus\fi
139  \ifx\SOUL@@,\else\SOUL@flushcomma\fi

140  \ifx\SOUL@@'\else\SOUL@flushapo\fi
141  \ifx\SOUL@@`\else\SOUL@flushgrave\fi
142  \ifx\SOUL@@-%
143  \advance\SOUL@minus\@ne
144  \else\ifx\SOUL@@,%
145  \advance\SOUL@comma\@ne
146  \else\ifx\SOUL@@'%
147  \advance\SOUL@apo\@ne
148  \else\ifx\SOUL@@`%
149  \advance\SOUL@grave\@ne
150  \else
151  \SOUL@flushminus
152  \SOUL@flushcomma
153  \SOUL@flushapo
154  \SOUL@flushgrave
155  \ifx\SOUL@@\SOUL@stop
156  \def\SOUL@n*{%
157  \SOUL@doword
158  \SOUL@eventuallyexhyphen\null

159 }%

160  \else\ifx\SOUL@@\par
161  \def\SOUL@n*\par{\par\leavevmode\SOUL@scan}%
162  \else\if\noexpand\SOUL@@\SOUL@spc
163  \SOUL@doword
164  \SOUL@eventuallyexhyphen\null
165  \ifSOUL@ignorespaces
166  \else
167  \SOUL@everyspace{}%
168  \fi
169  \def\SOUL@n* {\SOUL@scan}%
170  \else\ifx\SOUL@@\\%
171  \SOUL@doword
172  \SOUL@eventuallyexhyphen\null
173  \SOUL@everyspace{\unskip\nobreak\hfil\break}%
174  \SOUL@ignorespacestrue
175  \else\ifx\SOUL@@~%
176  \SOUL@doword
177  \SOUL@eventuallyexhyphen\null
178  \SOUL@everyspace{\nobreak}%
179  \else\ifx\SOUL@@\slash
180  \SOUL@doword
181  \SOUL@eventuallyexhyphen{/}%
182  \SOUL@exhyphen{/}%
183  \else\ifx\SOUL@@\mbox
184  \def\SOUL@n*{\SOUL@addprotect}%
185  \else\ifx\SOUL@@\hbox
186  \def\SOUL@n*{\SOUL@addprotect}%
187  \else\ifx\SOUL@@\soulomit
188  \def\SOUL@n*\soulomit##1{%
189  \SOUL@doword
190  {\spaceskip\SOUL@spaceskip##1}%
191  \SOUL@scan

192 }%

193  \else\ifx\SOUL@@\break

194  \SOUL@doword
195  \break
196  \else\ifx\SOUL@@\linebreak
197  \SOUL@doword
198  \SOUL@everyspace{\linebreak}%
199  \else\ifcat\bgroup\noexpand\SOUL@@
200  \def\SOUL@n*{\SOUL@addgroup{}}%
201  \else\ifcat$\noexpand\SOUL@@
202  \def\SOUL@n*{\SOUL@addmath}%
203  \else
204  \def\SOUL@n*{\SOUL@dotoken}%
205  \fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi
206  \fi\fi\fi\fi
207  \SOUL@n*%

208 }

\SOUL@flushminus \SOUL@flushcomma \SOUL@flushapo \SOUL@flushgrave As their names imply, these macros flush special tokens or token groups to the word register. They don’t do anything if the respective counter equals zero. \SOUL@minus does also flush the word register, because hyphens disturb the analyzer.

209 \def\SOUL@flushminus{%
210  \ifcase\SOUL@minus
211  \else
212  \SOUL@doword
213  \SOUL@eventuallyexhyphen{-}%
214  \ifcase\SOUL@minus
215  \or
216  \SOUL@exhyphen{-}%
217  \or
218  \SOUL@exhyphen{--}%
219  \or
220  \SOUL@exhyphen{---}%
221  \fi
222  \SOUL@minus\z@
223  \fi

224 }

225 \def\SOUL@flushcomma{%
226  \ifcase\SOUL@comma
227  \or
228  \edef\x{\SOUL@word={\the\SOUL@word,}}\x
229  \or
230  \edef\x{\SOUL@word={\the\SOUL@word{{,,}}}}\x
231  \fi
232  \SOUL@comma\z@

233 }

234 \def\SOUL@flushapo{%
235  \ifcase\SOUL@apo
236  \or
237  \edef\x{\SOUL@word={\the\SOUL@word'}}\x
238  \or
239  \edef\x{\SOUL@word={\the\SOUL@word{{''}}}}\x
240  \fi
241  \SOUL@apo\z@

242 }

243 \def\SOUL@flushgrave{%
244  \ifcase\SOUL@grave
245  \or
246  \edef\x{\SOUL@word={\the\SOUL@word`}}\x
247  \or
248  \edef\x{\SOUL@word={\the\SOUL@word{{``}}}}\x
249  \fi
250  \SOUL@grave\z@

251 }

\SOUL@dotoken Command sequences from the \SOUL@cmds list are handed over to \SOUL@docmd, everything else is added to \SOUL@word, which will be fed to the analyzer every time a word is completed. Since robust commands come with an additional space, we have also to examine if there’s a space variant. Otherwise we couldn’t detect pre-expanded formerly robust commands.

252 \def\SOUL@dotoken#1{%
253  \def\SOUL@@{\SOUL@addtoken{#1}}%
254  \def\\##1##2{%
255  \edef\SOUL@x{\string#1}%
256  \edef\SOUL@n{\string##2}%
257  \ifx\SOUL@x\SOUL@n
258  \def\SOUL@@{\SOUL@docmd{##1}{#1}}%
259  \else
260  \edef\SOUL@n{\string##2\space}%
261  \ifx\SOUL@x\SOUL@n
262  \def\SOUL@@{\SOUL@docmd{##1}{#1}}%
263  \fi
264  \fi

265 }%

266  \the\SOUL@cmds
267  \SOUL@@

268 }

\SOUL@docmd Here we deal with commands that were registered with \soulregister or \soulaccent or were already predefined in \SOUL@cmds. Commands with identifier 9 are accents that are put in a group with their argument. Identifier 8 is reserved for the \footnote command, and 7 for the \textsuperscript or similar commands. The others are mostly (but not necessarily) font switching commands, which may (1) or may not (0) take an argument. A registered command leads to the current word buffer being flushed to the analyzer, after which the command itself is executed.

Font switching commands which take an argument need special treatment: They need to increment the level counter, so that \SOUL@eval knows where to stop scanning. Furthermore the scanner has to be enabled to see the next token after the opening brace.

269 \def\SOUL@docmd#1#2{%
270  \ifx9#1%
271  \def\SOUL@@{\SOUL@addgroup{#2}}%
272  \else\ifx8#1%
273  \SOUL@doword
274  \def\SOUL@@##1{%
275  \SOUL@token={\footnotemark}%

276  \SOUL@everytoken
277  \SOUL@syllable={\footnotemark}%
278  \SOUL@everysyllable
279  \footnotetext{##1}%
280  \SOUL@doword
281  \SOUL@scan

282 }%

283  \else\ifx7#1%
284  \SOUL@doword
285  \def\SOUL@@##1{%
286  \SOUL@token={#2{##1}}%
287  \SOUL@everytoken
288  \SOUL@syllable={#2{##1}}%
289  \SOUL@everysyllable
290  \SOUL@doword
291  \SOUL@scan

292 }%

293  \else\ifx1#1%
294  \SOUL@doword
295  \def\SOUL@@##1{%
296  #2{\protect\SOUL@do{##1}}%
297  \SOUL@scan

298 }%

299  \else
300  \SOUL@doword
301  #2%
302  \let\SOUL@@\SOUL@scan
303  \fi\fi\fi\fi
304  \SOUL@@

305 }

\SOUL@addgroup \SOUL@addmath \SOUL@addprotect \SOUL@addtoken The macro names say it all. Each of these macros adds some token to the

word buffer \SOUL@word. Setting \protect is necessary to make things like

 \so{{a\itshape b}} work.
306 \def\SOUL@addgroup#1#2{%
307  {\let\protect\noexpand
308  \edef\x{\global\SOUL@word={\the\SOUL@word{{\noexpand#1#2}}}}\x}%
309  \SOUL@scan

310 }

311 \def\SOUL@addmath$#1${%
312  {\let\protect\noexpand
313  \edef\x{\global\SOUL@word={\the\SOUL@word{{\hbox{$#1$}}}}}\x}%
314  \SOUL@scan

315 }

316 \def\SOUL@addprotect#1#2{%
317  {\let\protect\noexpand
318  \edef\x{\global\SOUL@word={\the\SOUL@word{{\hbox{#2}}}}}\x}%
319  \SOUL@scan

320 }

321 \def\SOUL@addtoken#1{%
322  \edef\x{\SOUL@word={\the\SOUL@word\noexpand#1}}\x
323  \SOUL@scan

324 }

\SOUL@exhyphen Dealing with explicit hyphens can’t be done before we know the following character, because we need to know if a kerning value has to be inserted, hence we delay the \SOUL@everyexhyphen call. Unfortunately, the word scanner has no look-ahead mechanism.

325 \def\SOUL@exhyphen#1{%
326  \SOUL@getkern{\the\SOUL@lasttoken}{\SOUL@hyphkern}{#1}%
327  \gdef\SOUL@eventuallyexhyphen##1{%
328  \SOUL@getkern{#1}{\SOUL@charkern}{##1}%
329  \SOUL@everyexhyphen{#1}%
330  \gdef\SOUL@eventuallyexhyphen####1{}%

331 }% 332 }

\SOUL@cmds Here is a list of pre-registered commands that the analyzer cannot handle, so the scanner has to look after them. Every entry consists of a handle (\\), an identifier and the macro name. The class identifier can be 9 for accents, 8 for the \footnote command, 7 for the \textsuperscript command, 0 for commands without arguments and 1 for commands that take one argument. Commands with two or more arguments are not supported.

333 \SOUL@cmds={%
334  \\9\`\\9\'\\9\^\\9\"\\9\~\\9\=\\9\.%
335  \\9\u\\9\v\\9\H\\9\t\\9\c\\9\d\\9\b\\9\r
336  \\1\emph\\1\textrm\\1\textsf\\1\texttt\\1\textmd\\1\textbf
337  \\1\textup\\1\textsl\\1\textit\\1\textsc\\1\textnormal
338  \\0\rmfamily\\0\sffamily\\0\ttfamily\\0\mdseries\\0\upshape
339  \\0\slshape\\0\itshape\\0\scshape\\0\normalfont
340  \\0\em\\0\rm\\0\bf\\0\it\\0\tt\\0\sc\\0\sl\\0\sf
341  \\0\tiny\\0\scriptsize\\0\footnotesize\\0\small
342  \\0\normalsize\\0\large\\0\Large\\0\LARGE\\0\huge\\0\Huge
343  \\1\MakeUppercase\\7\textsuperscript\\8\footnote
344  \\1\textfrak\\1\textswab\\1\textgoth
345  \\0\frakfamily\\0\swabfamily\\0\gothfamily

346 }

\soulregister \soulfont \soulaccent Register a font switching command (or some other command) for the scanner. The first argument is the macro name, the second is the number of arguments

(0 or 1). Example: \soulregister{\bold}{0}. \soulaccent has only one

argument—the accent macro name. Example: \soulaccent{\~}. It is a shortcut for \soulregister{\~}{9}. The \soulfont command is a synonym for \soulregister and is kept for compatibility reasons.

347 \def\soulregister#1#2{{%
348  \edef\x{\global\SOUL@cmds={\the\SOUL@cmds
349  \noexpand\\#2\noexpand#1}}\x

350 }}

351 \def\soulaccent#1{\soulregister{#1}9}
352 \let\soulfont\soulregister

8.3 The analyzer

\SOUL@doword The only way to find out, where a given word can be broken into syllables, is to let TEX actually typeset the word under conditions that enforce every possible hyphenation. The result is a paragraph with one line for every syllable.

353 \def\SOUL@doword{%
354  \edef\x{\the\SOUL@word}%
355  \ifx\x\empty
356  \else
357  \SOUL@buffer={}%
358  \setbox\z@\vbox{%
359  \SOUL@tt
360  \hyphenchar\font`\-
361  \hfuzz\maxdimen
362  \hbadness\@M
363  \pretolerance\m@ne
364  \tolerance\@M
365  \leftskip\z@
366  \rightskip\z@
367  \hsize1sp
368  \everypar{}%
369  \parfillskip\z@\@plus1fil
370  \hyphenpenalty-\@M
371  \noindent
372  \hskip\z@
373  \relax
374  \the\SOUL@word}%
375  \let\SOUL@errmsg\SOUL@error
376  \let\-\relax
377  \count@\m@ne
378  \SOUL@analyze
379  \SOUL@word={}%
380  \fi

381 }

We store the hyphen width of the ectt1000 font, because we will need it in \SOUL@doword. (ectt1000 is a mono-spaced font, so every other character would have worked, too.)

382 \setbox\z@\hbox{\SOUL@tt-}
383 \newdimen\SOUL@ttwidth
384 \SOUL@ttwidth\wd\z@
385 \def\SOUL@sethyphenchar{%
386  \ifnum\hyphenchar\font=\m@ne
387  \else
388  \char\hyphenchar\font
389  \fi

390 }

\SOUL@analyze This macro decomposes the box that \SOUL@doword has built. Because we have to start at the bottom, we put every syllable onto the stack and execute ourselves recursively. If there are no syllables left, we return from the recursion and pick syllable after syllable from the stack again—this time from top to bottom—and hand the syllable width \SOUL@syllgoal over to \SOUL@dosyllable. All but the last syllable end with the hyphen character, hence we subtract the hyphen width accordingly. After processing a syllable we calculate the hyphen kern (i. e. the kerning amount between the last character and the hyphen). This might be needed by \SOUL@everyhyphen, which we call now.

391 \def\SOUL@analyze{{%

392  \setbox\z@\vbox{%
393  \unvcopy\z@
394  \unskip
395  \unpenalty
396  \global\setbox\@ne=\lastbox}%
397  \ifvoid\@ne
398  \else
399  \setbox\@ne\hbox{\unhbox\@ne}%
400  \SOUL@syllgoal=\wd\@ne
401  \advance\count@\@ne
402  \SOUL@analyze
403  \SOUL@syllwidth\z@
404  \SOUL@syllable={}%
405  \ifnum\count@>\z@
406  \advance\SOUL@syllgoal-\SOUL@ttwidth
407  \SOUL@dosyllable
408  \SOUL@getkern{\the\SOUL@lasttoken}{\SOUL@hyphkern}%
409  {\SOUL@sethyphenchar}%
410  \SOUL@everyhyphen
411  \else
412  \SOUL@dosyllable
413  \fi
414  \fi

415 }}

\SOUL@dosyllable This macro typesets token after token from \SOUL@word until \SOUL@syllwidth

has reached the requested width \SOUL@syllgoal. Furthermore the kerning values are prepared in case \SOUL@everytoken needs them. The \< command used by \so and \caps needs some special treatment: It has to be checked for, even before we can end a syllable.

416 \def\SOUL@dosyllable{%
417  \SOUL@gettoken
418  \SOUL@eventuallyexhyphen{\the\SOUL@token}%
419  \edef\x{\the\SOUL@token}%
420  \ifx\x\SOUL@hyphenhintM
421  \let\SOUL@n\SOUL@dosyllable
422  \else\ifx\x\SOUL@lowerthanM
423  \SOUL@gettoken
424  \SOUL@getkern{\the\SOUL@lasttoken}{\SOUL@charkern}
425  {\the\SOUL@token}%
426  \SOUL@everylowerthan
427  \SOUL@puttoken
428  \let\SOUL@n\SOUL@dosyllable
429  \else\ifdim\SOUL@syllwidth=\SOUL@syllgoal
430  \SOUL@everysyllable
431  \SOUL@puttoken
432  \let\SOUL@n\relax
433  \else\ifx\x\SOUL@stopM
434  \SOUL@errmsg
435  \global\let\SOUL@errmsg\relax
436  \let\SOUL@n\relax
437  \else
438  \setbox\tw@\hbox{\SOUL@tt\the\SOUL@token}%
439  \advance\SOUL@syllwidth\wd\tw@

440  \global\SOUL@lasttoken=\SOUL@token
441  \SOUL@gettoken
442  \SOUL@getkern{\the\SOUL@lasttoken}{\SOUL@charkern}
443  {\the\SOUL@token}%
444  \SOUL@puttoken
445  \global\SOUL@token=\SOUL@lasttoken
446  \SOUL@everytoken
447  \edef\x{\SOUL@syllable={\the\SOUL@syllable\the\SOUL@token}}\x
448  \let\SOUL@n\SOUL@dosyllable
449  \fi\fi\fi\fi
450  \SOUL@n

451 }

\SOUL@gettoken Provide the next token in \SOUL@token. If there’s already one in the buffer, use that one first.

452 \def\SOUL@gettoken{%
453  \edef\x{\the\SOUL@buffer}%
454  \ifx\x\empty
455  \SOUL@nexttoken
456  \else
457  \global\SOUL@token=\SOUL@buffer
458  \global\SOUL@buffer={}%
459  \fi

460 }

\SOUL@puttoken The possibility to put tokens back makes the scanner design much cleaner. There’s only room for one token, though, so we issue an error message if \SOUL@puttoken is told to put a token back while the buffer is still in use. Note that \SOUL@debug is actually undefined. This won’t hurt as it can only happen during driver design. No user will ever see this message.

461 \def\SOUL@puttoken{%
462  \edef\x{\the\SOUL@buffer}%
463  \ifx\x\empty
464  \global\SOUL@buffer=\SOUL@token
465  \global\SOUL@token={}%
466  \else
467  \SOUL@debug{puttoken called twice}%
468  \fi

469 }

\SOUL@nexttoken \SOUL@splittoken If the word buffer \SOUL@word is empty, deliver a \SOUL@stop, otherwise take the

next token.

470 \def\SOUL@nexttoken{%
471  \edef\x{\the\SOUL@word}%
472  \ifx\x\empty
473  \SOUL@token={\SOUL@stop}%
474  \else
475  \expandafter\SOUL@splittoken\the\SOUL@word\SOUL@stop
476  \fi

477 }

478 \def\SOUL@splittoken#1#2\SOUL@stop{%
479  \global\SOUL@token={#1}%
480  \global\SOUL@word={#2}%

\SOUL@getkern Assign the kerning value between the first and the third argument to the second,

 which has to be a \dimen register. \SOUL@getkern{A}{\dimen0}{V} will assign

the kerning value between ‘A’ and ‘V’ to \dimen0.

482 \def\SOUL@getkern#1#2#3{%
483  \setbox\tw@\hbox{#1#3}%
484  #2\wd\tw@
485  \setbox\tw@\hbox{#1\null#3}%
486  \advance#2-\wd\tw@

487 }

\SOUL@setkern Set a kerning value if it doesn’t equal 0 pt. Of course, we could also set a zero value, but that would needlessly clutter the logfile.

488 \def\SOUL@setkern#1{\ifdim#1=\z@\else\kern#1\fi}

\SOUL@error This error message will be shown once for every word that couldn’t be recon-

 structed by \SOUL@dosyllable.
489 \def\SOUL@error{%
490  \vrule\@height.8em\@depth.2em\@width1em
491  \PackageError{soul}{Reconstruction failed}{%
492  I came across hyphenatable material enclosed in group
493  braces,^^Jwhich I can't handle. Either drop the braces or
494  make the material^^Junbreakable using an \string\mbox\space
495  (\string\hbox). Note that a space^^Jalso counts as possible
496  hyphenation point. See page 4 of the manual.^^JI'm leaving
497  a black square so that you can see where I am right now.%

498 }% 499 }

\SOUL@setup This is a null driver, that will be used as the basis for other drivers. These have then to redefine only interface commands that shall differ from the default.

500 \def\SOUL@setup{%
501  \let\SOUL@preamble\relax
502  \let\SOUL@postamble\relax
503  \let\SOUL@everytoken\relax
504  \let\SOUL@everysyllable\relax
505  \def\SOUL@everyspace##1{##1\space}%
506  \let\SOUL@everyhyphen\relax
507  \def\SOUL@everyexhyphen##1{##1}%
508  \let\SOUL@everylowerthan\relax

509 }

510 \SOUL@setup

8.4 The l e t t e r s p a c i n g driver

\SOUL@sosetletterskip A handy helper macro that sets the inter-letter skip with a draconian \penalty.

511 \def\SOUL@sosetletterskip{\nobreak\hskip\SOUL@soletterskip}

\SOUL@sopreamble If letterspacing (\so or \caps) follows a white space, we replace it with our outer space. L^{A}TEX uses \hskip1sp as marker in tabular entries, so we ignore tiny skips.

512 \def\SOUL@sopreamble{%
513  \ifdim\lastskip>5sp
514  \unskip
515  \hskip\SOUL@soouterskip
516  \fi
517  \spaceskip\SOUL@soinnerskip

518 }

\SOUL@sopostamble Start the look-ahead scanner \SOUL@socheck outside the \SOUL@ scope. That’s why we make the outer space globally available in \skip@.

519 \def\SOUL@sopostamble{%
520  \global\skip@=\SOUL@soouterskip
521  \aftergroup\SOUL@socheck

522 }

\SOUL@socheck \SOUL@sodoouter Read the next token after the soul command into \SOUL@@ and examine it. If it’s some kind of space, replace it with outer space and the appropriate penalty, else if it’s a closing brace, continue scanning. If it is neither: do nothing.

523 \def\SOUL@socheck{%
524  \futurelet\SOUL@@\SOUL@sodoouter

525 }

526 \def\SOUL@sodoouter{%
527  \def\SOUL@n*##1{\hskip\skip@}%
528  \ifcat\egroup\noexpand\SOUL@@
529  \unkern
530  \egroup
531  \def\SOUL@n*{\afterassignment\SOUL@socheck\let\SOUL@x=}%
532  \else\ifx\SOUL@spc\SOUL@@
533  \def\SOUL@n* {\hskip\skip@}%
534  \else\ifx~\SOUL@@
535  \def\SOUL@n*~{\nobreak\hskip\skip@}%
536  \else\ifx\ \SOUL@@
537  \else\ifx\space\SOUL@@
538  \else\ifx\@xobeysp\SOUL@@
539  \else
540  \def\SOUL@n*{}%
541  \let\SOUL@@\relax
542  \fi\fi\fi\fi\fi\fi
543  \SOUL@n*%

544 }

\SOUL@soeverytoken Typeset the token and put an unbreakable inter-letter skip thereafter. If the token is \< then remove the last skip instead. Gets the character kerning value between the actual and the next token in \SOUL@charkern.

545 \def\SOUL@soeverytoken{%
546  \edef\x{\the\SOUL@token}%
547  \ifx\x\SOUL@lowerthanM
548  \else
549  \global\let\SOUL@soeventuallyskip\SOUL@sosetletterskip
550  \the\SOUL@token
551  \SOUL@gettoken
552  \edef\x{\the\SOUL@token}%
553  \ifx\x\SOUL@stopM

554  \else
555  \SOUL@setkern\SOUL@charkern
556  \SOUL@sosetletterskip
557  \SOUL@puttoken
558  \fi
559  \fi

560 }

\SOUL@soeveryspace This macro sets an inner space. The argument may contain penalties and is used for the ~ command. This construction was needed to make colored underlines work, without having to put any of the coloring commands into the core. \kern\z@ prevents in subsequent \so commands that the second discards the outer space of the first. To remove the space simply use \unkern\unskip.

561 \def\SOUL@soeveryspace#1{#1\space\kern\z@}

\SOUL@soeveryhyphen Sets implicit hyphens. The kerning value between the current token and the hyphen character is passed in \SOUL@hyphkern.

562 \def\SOUL@soeveryhyphen{%
563  \discretionary{%
564  \unkern
565  \SOUL@setkern\SOUL@hyphkern
566  \SOUL@sethyphenchar
567  }{}{}%

568 }

\SOUL@soeveryexhyphen Sets the explicit hyphen that is passed as argument. \SOUL@soeventuallyskip equals \SOUL@sosetletterskip, except when a \< had been detected. This is necessary because \SOUL@soeveryexhyphen wouldn’t know otherwise, that it follows

a \<.
569 \def\SOUL@soeveryexhyphen#1{%
570  \SOUL@setkern\SOUL@hyphkern
571  \SOUL@soeventuallyskip
572  \hbox{#1}%
573  \discretionary{}{}{%
574  \SOUL@setkern\SOUL@charkern

575 }%

576  \SOUL@sosetletterskip
577  \global\let\SOUL@soeventuallyskip\relax

578 }

\SOUL@soeverylowerthan Let \< remove the last inter-letter skip. Set the kerning value between the token before and that after the \< command.

579 \def\SOUL@soeverylowerthan{%
580  \unskip
581  \unpenalty
582  \global\let\SOUL@soeventuallyskip\relax
583  \SOUL@setkern\SOUL@charkern

584 }

\SOUL@sosetup Override all interface macros by our letterspacing versions. The only unused macro

is \SOUL@everysyllable.
585 \def\SOUL@sosetup{%

586  \SOUL@setup
587  \let\SOUL@preamble\SOUL@sopreamble
588  \let\SOUL@postamble\SOUL@sopostamble
589  \let\SOUL@everytoken\SOUL@soeverytoken
590  \let\SOUL@everyspace\SOUL@soeveryspace
591  \let\SOUL@everyhyphen\SOUL@soeveryhyphen
592  \let\SOUL@everyexhyphen\SOUL@soeveryexhyphen
593  \let\SOUL@everylowerthan\SOUL@soeverylowerthan

594 }

\SOUL@setso A handy macro for internal use.

595 \def\SOUL@setso#1#2#3{%
596  \def\SOUL@soletterskip{#1}%
597  \def\SOUL@soinnerskip{#2}%
598  \def\SOUL@soouterskip{#3}%

599 }

\sodef This macro assigns the letterspacing skips as well as an optional font switching command to a command sequence name. \so itself will be defined using this macro.

600 \def\sodef#1#2#3#4#5{%
601  \DeclareRobustCommand*#1{\SOUL@sosetup
602  \def\SOUL@preamble{%
603  \SOUL@setso{#3}{#4}{#5}%
604  #2%
605  \SOUL@sopreamble

606 }%

607  \SOUL@

608 }% 609 }

\resetso Let \resetso define reasonable default values for letterspacing.

610 \def\resetso{%
611  \sodef\textso{}{.25em}{.65em\@plus.08em\@minus.06em}%
612  {.55em\@plus.275em\@minus.183em}%

613 }

614 \resetso

\sloppyword Set up a letterspacing macro that inserts slightly stretchable space between the characters. This can be used to typeset long words in narrow columns, where ragged paragraphs are undesirable. See section 6.3.

615 \sodef\sloppyword{%
616  \linepenalty10
617  \hyphenpenalty10
618  \adjdemerits\z@
619  \doublehyphendemerits\z@
620  \finalhyphendemerits\z@
621  \emergencystretch.1em}%
622  {\z@\@plus.1em}%
623  {.33em\@plus.11em\@minus.11em}%
624  {.33em\@plus.11em\@minus.11em}

8.5 The caps driver

\caps Unless run under L^{A}TEX, make \caps just another simple letterspacing macro that selects a font \capsfont (defaulting to \relax) but doesn’t have any special capabilities.

625 \ifx\documentclass\@undefined
626 \let\capsfont\relax
627 \let\capsreset\relax
628 \def\capsdef#1#2#3#4#5{}
629 \def\capssave#1{}
630 \def\capsselect#1{}
631 \sodef\textcaps{\capsfont}
632  {.028em\@plus.005em\@minus.01em}%
633  {.37em\@plus.1667em\@minus.111em}%
634  {.37em\@plus.1em\@minus.14em}

\capsreset … else, if run under L^{A}TEX prepare a set of macros that maintain a database with certain letterspacing values for different fonts. \capsreset clears the database and inserts a default rule.

635 \else
636 \DeclareRobustCommand*\capsreset{%
637  \let\SOUL@capsbase\empty
638  \SOUL@capsdefault

639 }

\capsdef Add an entry to the database, which is of course nothing else than a TEX macro. See section “List macros” of appendix D in the TEXbook [5] for details.

640 \def\capsdef#1#2#3#4#5{{%
641  \toks\z@{\\{#1/#2/#3/#4/#5}}%
642  \toks\tw@=\expandafter{\SOUL@capsbase}%
643  \xdef\SOUL@capsbase{\the\toks\z@\the\toks\tw@}%

644 }}

\capssave \capsselect Save the current database in a macro within the SOUL@ namespace and let \capsselect restore this database.

645 \DeclareRobustCommand*\capssave[1]{%
646  \expandafter\global\expandafter\let
647  \csname SOUL@db@#1\endcsname\SOUL@capsbase

648 }

649 \DeclareRobustCommand*\capsselect[1]{%
650  \expandafter\let\expandafter\SOUL@capsbase
651  \csname SOUL@db@#1\endcsname

652 }

\SOUL@capsfind \SOUL@caps Go through the database entries and pick the first entry that matches the currently active font. Then define an internal macro that uses the respective spacing values in a macro that is equivalent to the \textso command.

653 \def\SOUL@capsfind#1/#2/#3/#4/#5/#6/#7/#8/#9/{%
654  \let\SOUL@match=1%
655  \SOUL@chk{#1}\f@encoding
656  \SOUL@chk{#2}\f@family
657  \SOUL@chk{#3}\f@series

658  \SOUL@chk{#4}\f@shape
659  \SOUL@dimchk{#5}\f@size
660  \if\SOUL@match1%
661  \let\\\@gobble
662  \gdef\SOUL@caps{%
663  \SOUL@sosetup
664  \def\SOUL@preamble{\SOUL@setso{#7}{#8}{#9}#6%
665  \SOUL@sopreamble}%
666  \SOUL@}%
667  \fi

668 }

\SOUL@chk Sets the \SOUL@match flag if both parameters are equal. This is used for all NFSS elements except the font size.

669 \def\SOUL@chk#1#2{%
670  \if$#1$%
671  \else
672  \def\SOUL@n{#1}%
673  \ifx#2\SOUL@n\else\let\SOUL@match=0\fi
674  \fi

675 }

\SOUL@dimchk \SOUL@rangechk We do not only want to check if a given font size #1 matches #2, but also if it fits into a given range. An omitted lower boundary is replaced by \z@ and an omitted upper boundary by \maxdimen. The first of a series of \SOUL@chk and \SOUL@dimchk statements, which detects that the arguments don’t match, sets the \SOUL@match flag to zero. A value of 1 indicates that an entry in the font database matches the currently used font.

676 \def\SOUL@dimchk#1#2{\if$#1$\else\SOUL@rangechk{#2}#1--\@ne\@@\fi}
677 \def\SOUL@rangechk#1#2-#3-#4\@@{%
678  \count@=#4%
679  \ifnum\count@>\z@
680  \ifdim#1\p@=#2\p@\else\let\SOUL@match=0\fi
681  \else
682  \dimen@=\if$#2$\z@\else#2\p@\fi
683  \ifdim#1\p@<\dimen@\let\SOUL@match=0\fi
684  \dimen@=\if$#3$\maxdimen\else#3\p@\fi
685  \ifdim#1\p@<\dimen@\else\let\SOUL@match=0\fi
686  \fi

687 }

\textcaps Find a matching entry in the database and start the letterspacing mechanism with the given spacing values.

688 \DeclareRobustCommand*\textcaps{{%
689  \def\\##1{\expandafter\SOUL@capsfind##1/}%
690  \SOUL@capsbase
691  \aftergroup\SOUL@caps

692 }}

\SOUL@capsdefault Define a default database entry and a default font.

693 \def\SOUL@capsdefault{%
694  \capsdef{////}%

695  \SOUL@capsdfltfnt
696  {.028em\@plus.005em\@minus.01em}%
697  {.37em\@plus.1667em\@minus.1em}%
698  {.37em\@plus.111em\@minus.14em}%

699 }

700 \let\SOUL@capsdfltfnt\scshape
701 \capsreset
702 \fi

8.6 The underlining driver

\SOUL@ulleaders This macro sets the underline under the following \hskip.

703 \newdimen\SOUL@uldp
704 \newdimen\SOUL@ulht
705 \def\SOUL@ulleaders{%
706  \leaders\hrule\@depth\SOUL@uldp\@height\SOUL@ulht\relax

707 }

\SOUL@ulunderline Set an underline under the given material. It draws the line first, and the given material afterwards. This is needed for highlighting, but gives less than optimal results for colored overstriking, which, however, will hardly ever be used, anyway.

708 \def\SOUL@ulunderline#1{{%
709  \setbox\z@\hbox{#1}%
710  \dimen@=\wd\z@
711  \dimen@i=\SOUL@uloverlap
712  \advance\dimen@2\dimen@i
713  \rlap{%
714  \null
715  \kern-\dimen@i
716  \SOUL@ulcolor{\SOUL@ulleaders\hskip\dimen@}%

717 }%

718  \unhcopy\z@

719 }}

\SOUL@ulpreamble Just set up the line dimensions and the space skip. Normally, \spaceskip is unset and not used by TEX. We need it, though, because we feed it to the \leaders primitive.

720 \def\SOUL@ulpreamble{%
721  \SOUL@uldp=\SOUL@uldepth
722  \SOUL@ulht=-\SOUL@uldp
723  \advance\SOUL@uldp\SOUL@ulthickness
724  \spaceskip\SOUL@spaceskip

725 }

\SOUL@uleverysyllable By using \SOUL@everysyllable we don’t have to care about kerning values and get better results for highlighting, where negative kerning values would otherwise cut off characters.

726 \def\SOUL@uleverysyllable{%
727  \SOUL@ulunderline{%
728  \the\SOUL@syllable
729  \SOUL@setkern\SOUL@charkern

730 }% 731 }

\SOUL@uleveryspace Set a given penalty and an underlined \space equivalent. The \null prevents a nasty gap in \textfrak {a \textswab{b}}, while it doesn’t seem to hurt in all other cases. I didn’t investigate this.

732 \def\SOUL@uleveryspace#1{%
733  \SOUL@ulcolor{%
734  #1%
735  \SOUL@ulleaders
736  \hskip\spaceskip

737 }%

738  \null

739 }

\SOUL@uleveryhyphen If hyphenation takes place, output an underlined hyphen with the required hyphen kerning value.

740 \def\SOUL@uleveryhyphen{%
741  \discretionary{%
742  \unkern
743  \SOUL@ulunderline{%
744  \SOUL@setkern\SOUL@hyphkern
745  \SOUL@sethyphenchar

746 }%

747  }{}{}%

748 }

\SOUL@uleveryexhyphen Underline the given hyphen, en-dash, em-dash or \slash and care for kerning.

749 \def\SOUL@uleveryexhyphen#1{%
750  \SOUL@setkern\SOUL@hyphkern
751  \SOUL@ulunderline{#1}%
752  \discretionary{}{}{%
753  \SOUL@setkern\SOUL@charkern

754 }% 755 }

\SOUL@ulcolor \setulcolor Define the underline color or turn off coloring, in which case the lines are not just colored black, but remain uncolored. This makes them appear black, nevertheless, and has the advantage, that no Postscript \specials are cluttering the output.

756 \let\SOUL@ulcolor\relax
757 \def\setulcolor#1{%
758  \if$#1$
759  \let\SOUL@ulcolor\relax
760  \else
761  \def\SOUL@ulcolor{\textcolor{#1}}%
762  \fi

763 }

\setuloverlap \SOUL@uloverlap Set the overlap amount, that helps to avoid gaps on sloppy output devices.

764 \def\setuloverlap#1{\def\SOUL@uloverlap{#1}}
765 \setuloverlap{.25pt}

\SOUL@ulsetup The underlining driver is quite simple. No need for \SOUL@postamble and

\SOUL@everytoken.
766 \def\SOUL@ulsetup{%

767  \SOUL@setup
768  \let\SOUL@preamble\SOUL@ulpreamble
769  \let\SOUL@everysyllable\SOUL@uleverysyllable
770  \let\SOUL@everyspace\SOUL@uleveryspace
771  \let\SOUL@everyhyphen\SOUL@uleveryhyphen
772  \let\SOUL@everyexhyphen\SOUL@uleveryexhyphen

773 }

\SOUL@textul Describing self-explanatory macros is so boring!

774 \DeclareRobustCommand*\textul{\SOUL@ulsetup\SOUL@}

\setul \SOUL@uldepth \SOUL@ulthickness Set the underlining dimensions. Either value may be omitted and lets the respective macro keep its current contents.

775 \def\setul#1#2{%
776  \if$#1$\else\def\SOUL@uldepth{#1}\fi
777  \if$#2$\else\def\SOUL@ulthickness{#2}\fi

778 }

\resetul Set reasonable default values that fit most latin fonts.

779 \def\resetul{\setul{.65ex}{.1ex}}
780 \resetul

\setuldepth This macro sets all designated “letters” (\catcode=11) or the given material in a box and sets the underlining dimensions according to the box depth.

781 \def\setuldepth#1{{%
782  \def\SOUL@n{#1}%
783  \setbox\z@\hbox{%
784  \tracinglostchars\z@
785  \ifx\SOUL@n\empty
786  \count@\z@
787  \loop
788  \ifnum\catcode\count@=11\char\count@\fi
789  \ifnum\count@<\@cclv
790  \advance\count@\@ne
791  \repeat
792  \else
793  #1%
794  \fi

795 }%

796  \dimen@\dp\z@
797  \advance\dimen@\p@
798  \xdef\SOUL@uldepth{\the\dimen@}%

799 }}

8.7 The overstriking driver

\SOUL@stpreamble Striking out is just underlining with a raised line of a different color. Hence we only need to define the preamble accordingly and let the underlining preamble finally do its job. Not that colored overstriking was especially useful, but we want at least to keep it black while we might want to set underlines in some fancy color.

800 \def\SOUL@stpreamble{%
801  \dimen@\SOUL@ulthickness

802  \dimen@i=-.5ex
803  \advance\dimen@i-.5\dimen@
804  \edef\SOUL@uldepth{\the\dimen@i}%
805  \let\SOUL@ulcolor\SOUL@stcolor
806  \SOUL@ulpreamble

807 }

\SOUL@stsetup We re-use the whole underlining setup and just replace the preamble with our modified version.

808 \def\SOUL@stsetup{%
809  \SOUL@ulsetup
810  \let\SOUL@preamble\SOUL@stpreamble

811 }

\textst These pretzels are making me thirsty …

812 \DeclareRobustCommand*\textst{\SOUL@stsetup\SOUL@}

\SOUL@stcolor \setstcolor Set the overstriking color. This won’t be used often, but is required in cases, where the underlines are colored. You wouldn’t want to overstrike in the same color. Note that overstriking lines are drawn beneath the text, hence bright colors won’t look good.

813 \let\SOUL@stcolor\relax
814 \def\setstcolor#1{%
815  \if$#1$
816  \let\SOUL@stcolor\relax
817  \else
818  \def\SOUL@stcolor{\textcolor{#1}}%
819  \fi

820 }

8.8 The highlighting driver

\SOUL@hlpreamble This is nothing else than overstriking with very thick lines.

821 \def\SOUL@hlpreamble{%
822  \setul{}{2.5ex}%
823  \let\SOUL@stcolor\SOUL@hlcolor
824  \SOUL@stpreamble

825 }

\SOUL@hlsetup No need to re-invent the wheel. Just use the overstriking setup with a different preamble.

826 \def\SOUL@hlsetup{%
827  \SOUL@stsetup
828  \let\SOUL@preamble\SOUL@hlpreamble

829 }

\texthl \sethlcolor \SOUL@hlcolor Define the highlighting macro and the color setting macro with a simple default color. Yellow isn’t really pleasing, but it’s already predefined by the color package.

830 \DeclareRobustCommand*\texthl{\SOUL@hlsetup\SOUL@}
831 \def\sethlcolor#1{\if$#1$\else\def\SOUL@hlcolor{\textcolor{#1}}\fi}
832 \sethlcolor{yellow}

The package postamble

\so \ul \st \hl \caps OK, I lied. The short macro names are just abbreviations for their longer counterpart. Some people might be used to \text* style commands to select a certain font. And then it doesn’t hurt to reserve these early enough.

833 \let\so\textso
834 \let\ul\textul
835 \let\st\textst
836 \let\hl\texthl
837 \let\caps\textcaps

When used in an environment other than L^{A}TEX and the german package was already loaded, define the double quotes as accent.

838 \ifx\documentclass\@undefined
839  \ifx\mdqoff\@undefined
840  \else
841  \soulaccent{"}%
842  \fi
843  \catcode`\@=\atcode

If we have been loaded by a L^{A}TEX environment and the color package wasn’t also loaded, we disable all colors. Then we add the umlaut accent " if the german package is present. The quotes character has to be \catcode’d \active now, or it won’t get recognized later. The capsdefault option overrides the \caps driver and lets \SOUL@ set an underline. Finally load the local configuration, process the capsdefault option and exit.

844 \else
845  \bgroup
846  \catcode`\"\active
847  \AtBeginDocument{%
848  \@ifundefined{color}{%
849  \let\SOUL@color\relax
850  \let\setulcolor\@gobble
851  \let\setstcolor\@gobble
852  \let\sethlcolor\@gobble
853  \let\hl\ul
854  }{}
855  \@ifundefined{mdqoff}{}{\soulaccent{"}}

856 }

857  \egroup
858  \DeclareOption{capsdefault}{%
859  \AtBeginDocument{%
860  \def\SOUL@capsdfltfnt#1{%
861  \SOUL@ulsetup
862  \SOUL@ulpreamble
863  \scshape

864 }% 865 }% 866 }

867  \InputIfFileExists{soul.cfg}%
868  {\PackageInfo{soul}{Local config file soul.cfg used}}{}
869  \ProcessOptions
870 \fi
871 \endinput

$Id: soul.dtx,v 1.128 2003/11/17 22:57:24 m Rel $

49